# Classifying localizing subcategories of a Grothendieck category

### v1

@@ 4. The atom spectrum of a locally coherent category
Since $\mathcal{A}$ is locally coherent, it is clear that for every $\alpha\in\operatorname{ASpec}\mathcal{A}$, there exists a finitely presented object $M$ in $\mathcal{A}$ such that $\alpha\in\operatorname{ASupp}M$. If $M_{1}$ and $M_{2}$ are finitely presented objects in $\mathcal{A}$ and $\alpha\in\operatorname{ASupp}M_{1}\cap\operatorname{ASupp}M_{2}$, then by section 3, there exists a nonzero morphism $f_{i}:M_{i}\rightarrow E(\alpha)$ for $i=1,2$. Since $E(\alpha)$ is uniform, $\operatorname{Im}f_{1}\cap\operatorname{Im}f_{2}$ is a nonzero subobject of $E(\alpha)$ and so it contains a nonzero finitely generated subobject $X$ as $\mathcal{A}$ is locally coherent. Using the pull-back diagram and section 4, there exists a finitely presented subobject $L_{i}$ of $M_{i}$ such that $f_{i}(L_{i})=X$ for $i=1,2$. This implies that $X\in\overset{\longrightarrow}{\surd{M_{1}}}\cap\overset{\longrightarrow}{\surd{M_{2}}}$. By virtue of [H, Proposition 2.3], the morphism $f_{1}:L_{1}\rightarrow X$ factors through a quotient $N$ of $L_{1}$ which lies in $\surd{M_{2}}$. Therefore $N$ is finitely presented and $\alpha\in\operatorname{ASupp}N\subset\operatorname{ASupp}M_{1}\cap\operatorname{ASupp}M_{2}$. To prove the second claim, every nonzero finitely presented object $M$ in $\mathcal{A}$ has a simple quotient object $S$ so that $\alpha=\overline{S}\in\operatorname{ASupp}M$. ∎
To prevent any misunderstanding, we use the symbol $\operatorname{ZASpec}\mathcal{A}$ instead of $\operatorname{ASpec}\mathcal{A}$ with the new topology based on section 4. We notice that m-$\operatorname{ASpec}\mathcal{A}$ is a dense subset of $\operatorname{ZASpec}\mathcal{A}$. Because if $M$ is a finitely presented object in $\mathcal{A}$, it contains a maximal subobject $N$ so that the maximal atom $\overline{M/N}\in\operatorname{ASupp}M$. We also remark that if $M$ is a finitely presented object in $\mathcal{A}$, then $\operatorname{ASupp}K$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$ for any subobject $K$ of $M$. Because $K=\bigcup_{i}K_{i}$ is the direct union of its finitely generated subobjects. Since $M$ is finitely presented, each $K_{i}$ is finitely presented so that $\operatorname{ASupp}K=\bigcup_{i}\operatorname{ASupp}K_{i}$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$.
[…]
Let $\mathcal{X}$ be a Serre subcategory of $\operatorname{fp-}\mathcal{A}$. Then the inclusion functor $\overrightarrow{\mathcal{X}}^{\bot}\subset\mathcal{A}$ establishes a homeomorphism $G:\operatorname{ZASpec}(\mathcal{A}/\overrightarrow{\mathcal{X}})\rightarrow\operatorname{ZASpec}\mathcal{A}\setminus\operatorname{ASupp}\mathcal{X}$ endowed with relative subspace topology. Furthermore, $G(\operatorname{ASupp}M_{\mathcal{X}})=\operatorname{ASupp}M\setminus\operatorname{ASupp}\mathcal{X}$ for $M\in\operatorname{fp-}\mathcal{A}$.
Proof.
See [K, Theorem 5.4]. ∎
[…]
@@
Assume that $\mathcal{X}$ is a Serre subcategory of $\operatorname{fp-}\mathcal{A}$ and $\mathcal{U}$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$. It is clear that $\operatorname{ASupp}\mathcal{X}$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$ and $\operatorname{ASupp}^{-1}\mathcal{U}\cap\operatorname{fp-}\mathcal{A}$ is a Serre subcategory of $\operatorname{fp-}\mathcal{A}$. In order to prove $\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})\cap\operatorname{fp-}\mathcal{A}=\mathcal{X}$, it suffices to show that $\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})\cap\operatorname{fp-}\mathcal{A}\subset\mathcal{X}$. Given $M\in\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})\cap\operatorname{fp-}\mathcal{A}$, we have $\operatorname{ASupp}M\subset\operatorname{ASupp}\mathcal{X}$ and so $\operatorname{ASupp}M\subset\operatorname{ASupp}\overset{\longrightarrow}{\mathcal{X}}$. Then $\operatorname{ASupp}M_{\mathcal{X}}=\operatorname{ASupp}M\setminus\operatorname{ASupp}\mathcal{X}=\emptyset$ so that $M_{\mathcal{X}}=0$ by section 4 and section 4. Hence $M\in\overset{\longrightarrow}{\mathcal{X}}$ and so $M=\underset{\rightarrow}{\rm lim}M_{i}$ is the direct limit of objects $M_{i}$ in $\mathcal{X}$. Since $M$ is finitely presented, it is a direct summand of some $M_{i}$ so that $M\in\mathcal{X}$. The fact that $\operatorname{ASupp}(\operatorname{ASupp}^{-1}\mathcal{U})=\mathcal{U}$ is straightforward. ∎
@@
Given an open subset $\mathcal{U}$ of $\operatorname{ZASpec}\mathcal{A}$, by [Kr, Theorem 2.8], $\overrightarrow{\operatorname{ASupp}^{-1}\mathcal{U}\cap\operatorname{fp-}\mathcal{A}}$ is a localizing subcategory of finite type of $\mathcal{A}$. It is easy to see that $\operatorname{ASupp}(\overrightarrow{\operatorname{ASupp}^{-1}\mathcal{U}\cap\operatorname{fp-}\mathcal{A}})=\mathcal{U}$. Also, for every localizing subcategory $\mathcal{X}$ of finite type of $\mathcal{A}$, it follows from section 4 and [Kr, Lemma 2.3] that
@@
If $\operatorname{ASupp}\mathcal{X}$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$, then section 4 implies that $\mathcal{X}=\overrightarrow{\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})\cap\operatorname{fp-}\mathcal{A}}$ is of finite type of $\mathcal{A}$. The converse is clear. ∎
@@
Hence section 4 implies that $M\in\overrightarrow{\sqrt{\langle\operatorname{ASupp}^{-1}\varPhi_{\lambda}\cap\operatorname{fp-}\mathcal{A}|\hskip 2.84544pt\lambda\in\varLambda\rangle}}$. Since $M$ is finitely presented, it is the direct summand of some object in $\sqrt{\langle\operatorname{ASupp}^{-1}\varPhi_{\lambda}\cap\operatorname{fp-}\mathcal{A}|\hskip 2.84544pt\lambda\in\varLambda\rangle}$ and so $M\in\sqrt{\langle\operatorname{ASupp}^{-1}\varPhi_{\lambda}\cap\operatorname{fp-}\mathcal{A}|\hskip 2.84544pt\lambda\in\varLambda\rangle}$. Therefore $M$ has a finite filtration of finitely presented subobjects $0=M_{0}\subset M_{1}\subset M_{2}\subset\dots\subset M_{n}=M$ such that $M_{i}/M_{i-1}$ is a subquotient of some finitely presented object $N_{\lambda_{i}}\in\operatorname{ASupp}^{-1}\varPhi_{\lambda_{i}}$ for $1\leq i\leq n$. Therefore $\operatorname{ASupp}M=\bigcup_{i=1}^{n}\varPhi_{\lambda_{i}}$
@@
Let $H$ be a monoform object in $\operatorname{fp-}\mathcal{A}$. If $H$ is not monoform object in $\mathcal{A}$, then there exists a subobject $N$ of $H$ and a nonzero finitely presented subobject $X$ of $H/N$ isomorphic to a subobject of $H$. Since $\operatorname{fp-}\mathcal{A}$ is abelian, using a pullback diagram and section 4, there exist finitely generated (and so finitely presented) subobjects $C\subset D$ of $H$ such that $X\cong D/C$. This implies that $H$ and $H/C$ has a common nonzero subobject which contradicts the fact that $H$ is a monoform object of $\operatorname{fp-}\mathcal{A}$. ∎
[…]
The map $\alpha_{f}\stackrel{{\scriptstyle g}}{{\mapsto}}\alpha$ embeds $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ as a subspace of $\operatorname{ZASpec}\mathcal{A}$. Moreover, if $\mathcal{A}$ is locally monoform, then $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ is dense in $\operatorname{ZASpec}\mathcal{A}$.
@@
Given $\alpha_{f}\in\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$, there exists a monoform object $H$ of the abelian category $\operatorname{fp-}\mathcal{A}$ such that $\alpha_{f}=\overline{H}$. Thus $\alpha=\overline{H}\in\operatorname{ZASpec}\mathcal{A}$ and so identifying $\alpha_{f}$ with $\alpha$, we show that $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ is a topological subspace of $\operatorname{ZASpec}\mathcal{A}$. It is clear that $g$ is injective and we show that $g$ is continuous. For any finitely presented object $M$ in $\mathcal{A}$, we prove that $g^{-1}(\operatorname{ASupp}M)=\operatorname{fASupp}M$. The above argument indicates that $\operatorname{fASupp}M\subset g^{-1}(\operatorname{ASupp}M)$. To prove the converse, assume that $\alpha_{f}\in g^{-1}(\operatorname{ASupp}M)$. By section 4, there exists a finitely presented monoform object $H$ in $\mathcal{A}$ such that $\alpha=\overline{H}$ and $H$ is a subquotient of $M$. Using section 4, we can choose such $H$ such that $H=L/K$, where $L$ is a finitely presented subobject of $M$. Since $\operatorname{fp-}\mathcal{A}$ is abelian, we deduce that $K$ is finitely presented; and hence $\alpha_{f}\in\operatorname{fASupp}M$. To prove the second claim, the assumption and section 4 imply that every finitely presented object $M$ contains a finitely presented monoform subobject $H$ in $\mathcal{A}$. Then $\overline{H}=\alpha_{f}\in\operatorname{ASupp}M\cap\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$. ∎
In view of section 4, for any object $M$ in $\mathcal{A}$, we define the following subsets of $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$:
According to section 4, $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ can be considered as a subspace of $\operatorname{ZASpec}\mathcal{A}$. Furthermore, we have the following corollary.
[…]
Let $M$ be a finitely presented object in $\mathcal{A}$. Then $\operatorname{fAAss}M=\operatorname{AAss}M$.
@@
The first claim follows easily. For the second claim, if $\alpha\in\operatorname{AAss}M$, then $M$ contains a monoform subobject $H$ in $\mathcal{A}$ such that $\alpha=\overline{H}$. Since $\mathcal{A}$ is locally coherent, we may assume that $H$ is finitely generated. Hence $H$ is finitely presented because $M$ is finitely presented. Moreover, it is clear that $H$ is a monoform object of $\operatorname{fp-}\mathcal{A}$ and consequently $\alpha_{f}\in\operatorname{fAAss}M$. To prove the second claim, if $\alpha_{f}\in\operatorname{fAAss}M$, then $M$ contains a monoform subobject $H$ of $\operatorname{fp-}\mathcal{A}$ such that $\alpha=\overline{H}$. It follows from section 4 that $H$ is a monoform object in $\mathcal{A}$ so that $\alpha\in\operatorname{AAss}M$. Conversely if $\alpha\in\operatorname{AAss}M$, since $M$ is finitely presented, it contains a monoform subobject $H$ such that $\overline{H}=\alpha$. ∎
@@
[…]
If $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}=\operatorname{ASpec}\mathcal{A}$ (i.e. if g in section 4 is surjective), then $\mathcal{A}$ is locally noetherian.
@@
has a maximal element $K$ by Zorn’s lemma. We claim that $M/K$ is a monoform object in $\mathcal{A}$. Otherwise, there exists a subobject $L$ of $M$ containing $K$ such that $M/K$ and $M/L$ has a common nonzero subobject $X$. The maximality $K$ implies that $L$ is finitely generated so that $M/L$ is finitely presented. Then we may assume that $X$ is finitely presented. Hence there exists a suboobject $K_{1}$ of $M$ containig $K$ such that $X\cong K_{1}/K$ is finitely presented. Again maximality implies that $K_{1}$ is finitely generated and so is finitely presented. Since $\operatorname{fp-}\mathcal{A}$ is abelian, we deduce that $K$ is finitely presented which is a contradiction. Therefore $\overline{M/K}\in\operatorname{ASpec}\mathcal{A}$ and so the assumption implies that $M/K$ contains a finitely presented monoform subobject $Y/K$. Again by maximality of $K$, the subobject $Y$ is finitely presented so that $K$ is finitely presented which is a contradiction. ∎
@@
By [Kr, Theorem 2.8], it suffices to show that $\mathcal{X}=\overrightarrow{\mathcal{X}\cap\operatorname{fp-}\mathcal{A}}$. The inclusion $\overrightarrow{\mathcal{X}\cap\operatorname{fp-}\mathcal{A}}\subset\mathcal{X}$ is clear. For the other side, since $\mathcal{A}$ is a locally coherent, $M=\bigcup_{i}M_{i}$ is the direct union of its finitely generated. Then it suffices to show that each $M_{i}$ is finitely presented. Hence, we may assume that $M$ is a finitely generated subobject in $\mathcal{X}$ and we have to prove that $M$ is finitely presented. Since $\mathcal{A}$ is locally coherent, there exists a finitely presented object $N$ in $\mathcal{A}$ and a subobject $L$ of $N$ such that $M=N/L$. If $L$ is not finitely generated, the set
[…]
has a maximal element $K$ by Zorn’s lemma. By a similar proof as given in section 4, we can deduce that $N/K$ is a monoform object in $\mathcal{A}$. Thus $\overline{N/K}\in\operatorname{ASupp}M$ and the assumption implies that $N/K$ contains a finitely presented subobject $Y$. Therefore by the maximality of $K$, $N$ has a finitely presented subobject $Z$ containing $K$ such that $Y=Z/K$. Since $\operatorname{fp-}\mathcal{A}$ is abelian, we conclude that $K$ is finitely presented which is a contradiction. Hence $L$ is finitely generated so that $M$ is finitely presented. ∎

### v2

@@ 3. The atom spectrum of a locally coherent category
Since $\mathcal{A}$ is locally coherent, it is clear that for every $\alpha\in\operatorname{ASpec}\mathcal{A}$, there exists a finitely presented object $M$ in $\mathcal{A}$ such that $\alpha\in\operatorname{ASupp}M$. If $M_{1}$ and $M_{2}$ are finitely presented objects in $\mathcal{A}$ and $\alpha\in\operatorname{ASupp}M_{1}\cap\operatorname{ASupp}M_{2}$, then by section˜3, there exists a nonzero morphism $f_{i}:M_{i}\rightarrow E(\alpha)$ for $i=1,2$. Since $E(\alpha)$ is uniform, $\operatorname{Im}f_{1}\cap\operatorname{Im}f_{2}$ is a nonzero subobject of $E(\alpha)$ and so it contains a nonzero finitely generated subobject $X$ as $\mathcal{A}$ is locally coherent. Using the pull-back diagram and section˜3, there exists a finitely presented subobject $L_{i}$ of $M_{i}$ such that $f_{i}(L_{i})=X$ for $i=1,2$. This implies that $X\in\overset{\longrightarrow}{\surd{M_{1}}}\cap\overset{\longrightarrow}{\surd{M_{2}}}$. By virtue of [H, Proposition 2.3], the morphism $f_{1}:L_{1}\rightarrow X$ factors through a quotient $N$ of $L_{1}$ which lies in $\surd{M_{2}}$. Therefore $N$ is finitely presented and $\alpha\in\operatorname{ASupp}N\subset\operatorname{ASupp}M_{1}\cap\operatorname{ASupp}M_{2}$. To prove the second claim, every nonzero finitely presented object $M$ in $\mathcal{A}$ has a simple quotient object $S$ so that $\alpha=\overline{S}\in\operatorname{ASupp}M$. ∎
To prevent any misunderstanding, we use the symbol $\operatorname{ZASpec}\mathcal{A}$ instead of $\operatorname{ASpec}\mathcal{A}$ with the new topology based on section˜3. We notice that m-$\operatorname{ASpec}\mathcal{A}$ is a dense subset of $\operatorname{ZASpec}\mathcal{A}$. Because if $M$ is a finitely presented object in $\mathcal{A}$, it contains a maximal subobject $N$ so that the maximal atom $\overline{M/N}\in\operatorname{ASupp}M$. We also remark that if $M$ is a finitely presented object in $\mathcal{A}$, then $\operatorname{ASupp}K$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$ for any subobject $K$ of $M$. Because $K=\bigcup_{i}K_{i}$ is the direct union of its finitely generated subobjects. Since $M$ is finitely presented, each $K_{i}$ is finitely presented so that $\operatorname{ASupp}K=\bigcup_{i}\operatorname{ASupp}K_{i}$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$.
Theorem 3.4.
@@
Assume that $\mathcal{X}$ is a Serre subcategory of $\operatorname{fp-}\mathcal{A}$ and $\mathcal{U}$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$. It is clear that $\operatorname{ASupp}\mathcal{X}$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$ and $\operatorname{ASupp}^{-1}\mathcal{U}\cap\operatorname{fp-}\mathcal{A}$ is a Serre subcategory of $\operatorname{fp-}\mathcal{A}$. In order to prove $\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})\cap\operatorname{fp-}\mathcal{A}=\mathcal{X}$, it suffices to show that $\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})\cap\operatorname{fp-}\mathcal{A}\subset\mathcal{X}$. Given $M\in\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})\cap\operatorname{fp-}\mathcal{A}$, we have $\operatorname{ASupp}M\subset\operatorname{ASupp}\mathcal{X}$ and so $\operatorname{ASupp}M\subset\operatorname{ASupp}\overset{\longrightarrow}{\mathcal{X}}$. It follows from [H, Proposition 2.15] that $F(M)$ is a finitely presented object in $\mathcal{A}/\overset{\longrightarrow}{\mathcal{X}}$, where $F:\mathcal{A}\rightarrow\mathcal{A}/\overset{\longrightarrow}{\mathcal{X}}$ is the canonical functor. Moreover, $\operatorname{ASupp}F(M)=\operatorname{ASupp}M\setminus\operatorname{ASupp}\mathcal{X}=\emptyset$ by [K, Proposition 5.6]; and hence $F(M)=0$ by section˜3. Therefore, $M\in\overset{\longrightarrow}{\mathcal{X}}$ and so $M=\underset{\rightarrow}{\rm lim}M_{i}$ is the direct limit of objects $M_{i}$ in $\mathcal{X}$. Since $M$ is finitely presented, it is a direct summand of some $M_{i}$ so that $M\in\mathcal{X}$. The fact that $\operatorname{ASupp}(\operatorname{ASupp}^{-1}\mathcal{U})=\mathcal{U}$ is straightforward. ∎
@@
Given an open subset $\mathcal{U}$ of $\operatorname{ZASpec}\mathcal{A}$, by [Kr, Theorem 2.8], $\overrightarrow{\operatorname{ASupp}^{-1}\mathcal{U}\cap\operatorname{fp-}\mathcal{A}}$ is a localizing subcategory of finite type of $\mathcal{A}$. It is easy to see that $\operatorname{ASupp}(\overrightarrow{\operatorname{ASupp}^{-1}\mathcal{U}\cap\operatorname{fp-}\mathcal{A}})=\mathcal{U}$. Also, for every localizing subcategory $\mathcal{X}$ of finite type of $\mathcal{A}$, it follows from section˜3 and [Kr, Lemma 2.3] that
@@
If $\operatorname{ASupp}\mathcal{X}$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$, then section˜3 implies that $\mathcal{X}=\overrightarrow{\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})\cap\operatorname{fp-}\mathcal{A}}$ is of finite type of $\mathcal{A}$. The converse is clear. ∎
@@
Hence section˜3 implies that $M\in\overrightarrow{\sqrt{\langle\operatorname{ASupp}^{-1}\varPhi_{\lambda}\cap\operatorname{fp-}\mathcal{A}|\hskip 2.84544pt\lambda\in\varLambda\rangle}}$. Since $M$ is finitely presented, it is the direct summand of some object in $\sqrt{\langle\operatorname{ASupp}^{-1}\varPhi_{\lambda}\cap\operatorname{fp-}\mathcal{A}|\hskip 2.84544pt\lambda\in\varLambda\rangle}$ and so $M\in\sqrt{\langle\operatorname{ASupp}^{-1}\varPhi_{\lambda}\cap\operatorname{fp-}\mathcal{A}|\hskip 2.84544pt\lambda\in\varLambda\rangle}$. Therefore $M$ has a finite filtration of finitely presented subobjects $0=M_{0}\subset M_{1}\subset M_{2}\subset\dots\subset M_{n}=M$ such that $M_{i}/M_{i-1}$ is a subquotient of some finitely presented object $N_{\lambda_{i}}\in\operatorname{ASupp}^{-1}\varPhi_{\lambda_{i}}$ for $1\leq i\leq n$. Therefore $\operatorname{ASupp}M=\bigcup_{i=1}^{n}\varPhi_{\lambda_{i}}$
@@
Let $H$ be a monoform object in $\operatorname{fp-}\mathcal{A}$. If $H$ is not monoform object in $\mathcal{A}$, then there exists a subobject $N$ of $H$ and a nonzero finitely presented subobject $X$ of $H/N$ isomorphic to a subobject of $H$. Since $\operatorname{fp-}\mathcal{A}$ is abelian, using a pullback diagram and section˜3, there exist finitely generated (and so finitely presented) subobjects $C\subset D$ of $H$ such that $X\cong D/C$. This implies that $H$ and $H/C$ has a common nonzero subobject which contradicts the fact that $H$ is a monoform object of $\operatorname{fp-}\mathcal{A}$. ∎
Corollary 3.9.
Let $M$ be a finitely presented object in $\mathcal{A}$. Then $\operatorname{fAAss}M=(\operatorname{AAss}M)_{f}$.
@@
If $\alpha\in\operatorname{AAss}M$, then $M$ contains a monoform subobject $H$ in $\mathcal{A}$ such that $\alpha=\overline{H}$. Since $\mathcal{A}$ is locally coherent, we may assume that $H$ is finitely generated. Hence $H$ is finitely presented because $M$ is finitely presented. Moreover, it is clear that $H$ is a monoform object of $\operatorname{fp-}\mathcal{A}$ and consequently $\alpha_{f}\in\operatorname{fAAss}M$. To prove the other side, if $\alpha_{f}\in\operatorname{fAAss}M$, then $M$ contains a monoform subobject $H$ of $\operatorname{fp-}\mathcal{A}$ such that $\alpha=\overline{H}$. It follows from section˜3 that $H$ is a monoform object in $\mathcal{A}$ so that $\alpha\in\operatorname{AAss}M$. Conversely if $\alpha\in\operatorname{AAss}M$, since $M$ is finitely presented, it contains a monoform subobject $H$ such that $\overline{H}=\alpha$. ∎
The atom spectrum $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ can be considered as a subspace of $\operatorname{ZASpec}\mathcal{A}$.
Proposition 3.10.
The map $\alpha_{f}\stackrel{{\scriptstyle g}}{{\mapsto}}\alpha$ embeds $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ as a subspace of $\operatorname{ZASpec}\mathcal{A}$. Moreover, if $\operatorname{fp-}\mathcal{A}$ is locally monoform, then $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ is dense in $\operatorname{ZASpec}\mathcal{A}$.
@@
Given $\alpha_{f}\in\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$, there exists a monoform object $H$ of the abelian category $\operatorname{fp-}\mathcal{A}$ such that $\alpha_{f}=\overline{H}$. Thus $\alpha=\overline{H}\in\operatorname{ZASpec}\mathcal{A}$ and so identifying $\alpha_{f}$ with $\alpha$, we show that $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ is a topological subspace of $\operatorname{ZASpec}\mathcal{A}$. It is clear that $g$ is injective and we show that $g$ is continuous. For any finitely presented object $M$ in $\mathcal{A}$, we prove that $g^{-1}(\operatorname{ASupp}M)=\operatorname{fASupp}M$. The above argument indicates that $\operatorname{fASupp}M\subset g^{-1}(\operatorname{ASupp}M)$. To prove the converse, assume that $\alpha_{f}\in g^{-1}(\operatorname{ASupp}M)$. By section˜3, there exists a finitely presented monoform object $H$ in $\mathcal{A}$ such that $\alpha=\overline{H}$ and $H$ is a subquotient of $M$. Using section˜3, we can choose such $H$ such that $H=L/K$, where $L$ is a finitely presented subobject of $M$. Since $\operatorname{fp-}\mathcal{A}$ is abelian, we deduce that $K$ is finitely presented; and hence $\alpha_{f}\in\operatorname{fASupp}M$. To prove the second claim, the assumption and section˜3 imply that every finitely presented object $M$ contains a finitely presented monoform subobject $H$ in $\mathcal{A}$. Then $\overline{H}=\alpha_{f}\in\operatorname{ASupp}M\cap\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$. ∎
@@
Theorem 3.11.
If $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}=\operatorname{ASpec}\mathcal{A}$ (i.e. if g in section˜3 is surjective), then $\mathcal{A}$ is locally noetherian.
@@
has a maximal element $K$ by Zorn’s lemma. We claim that $M/K$ is a monoform object in $\mathcal{A}$. Otherwise, there exists a subobject $L$ of $M$ containing $K$ such that $M/K$ and $M/L$ has a common nonzero subobject $X$. The maximality $K$ implies that $L$ is finitely generated so that $M/L$ is finitely presented. Then we may assume that $X$ is finitely presented. Hence there exists a subobject $K_{1}$ of $M$ containig $K$ such that $X\cong K_{1}/K$ is finitely presented. Again maximality implies that $K_{1}$ is finitely generated and so is finitely presented. Since $\operatorname{fp-}\mathcal{A}$ is abelian, we deduce that $K$ is finitely presented which is a contradiction. Therefore $\overline{M/K}\in\operatorname{ASpec}\mathcal{A}$ and so the assumption implies that $M/K$ contains a finitely presented monoform subobject $Y/K$. Again by maximality of $K$, the subobject $Y$ is finitely presented so that $K$ is finitely presented which is a contradiction. ∎
In view of section˜3, for any object $M$ in $\mathcal{A}$, we define the following subsets of $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$:
@@
By [Kr, Theorem 2.8], it suffices to show that $\mathcal{X}=\overrightarrow{\mathcal{X}\cap\operatorname{fp-}\mathcal{A}}$. The inclusion $\overrightarrow{\mathcal{X}\cap\operatorname{fp-}\mathcal{A}}\subset\mathcal{X}$ is clear. For the other side, since $\mathcal{A}$ is a locally coherent, $M=\bigcup_{i}M_{i}$ is the direct union of its finitely generated subobjects. Then it suffices to show that each $M_{i}$ is finitely presented. Hence, we may assume that $M$ is a finitely generated object in $\mathcal{X}$ and we have to prove that $M$ is finitely presented. Since $\mathcal{A}$ is locally coherent, there exists a finitely presented object $N$ in $\mathcal{A}$ and a subobject $L$ of $N$ such that $M=N/L$. If $L$ is not finitely generated, the set
$\{X\subset N|\hskip 2.84544ptL\subset X\hskip 2.84544pt{\rm and}\hskip 2.84544ptX\hskip 2.84544ptis\hskip 2.84544pt{\rm not}\hskip 2.84544pt{\rm finitely}\hskip 2.84544pt{\rm generated}\}$
has a maximal element $K$ by Zorn’s lemma. By a similar proof as given in section˜3, we can deduce that $N/K$ is a monoform object in $\mathcal{A}$. Thus $\overline{N/K}\in\operatorname{ASupp}M$ and the assumption implies that $N/K$ contains a finitely presented subobject $Y$. Therefore by the maximality of $K$, NN has a finitely presented subobject $Z$ containing $K$ such that $Y=Z/K$. Since $\operatorname{fp-}\mathcal{A}$ is abelian, we conclude that $K$ is finitely presented which is a contradiction. Hence $L$ is finitely generated so that $M$ is finitely presented. ∎